### v1

# Conditional Autoencoder for Generating BNS Waveforms with Tidal and Precession Effects

Abstract
Gravitational waves from binary neutron star mergers provide insights into dense matter physics and strong-field gravity, but waveform modeling remains computationally challenging. We develop a deep generative model for gravitational waveforms from BNS mergers, covering the late inspiral, merger, and ringdown, incorporating precession and tidal effects. Using the conditional autoencoder, our model efficiently generates waveforms with high fidelity across a broad parameter space, including component masses $(m_{1},m_{2})$, spin components $(S_{1x},S_{1y},S_{1z},S_{2x},S_{2y},S_{2z})$ and tidal deformability $(\Lambda_{1},\Lambda_{2})$. Trained on $3\times 10^{5}$ waveforms from the IMRPhenomXP_NRTidalv2 waveform model, it achieves an average overlap accuracy of 99.6% on the test set. The model significantly accelerates waveform generation. For a single sample, it requires $0.12$ seconds (s), compared to $0.38$ s for IMRPhenomXP_NRTidalv2 and $0.62$ s for IMRPhenomPv2_NRTidal, making it approximately 3 to 5 times faster. When generating $10^{3}$ waveforms, the network completes the task in $0.86$ s, while traditional waveform approximation methods take over 4646464653535353 s. Our model achieves a total time of $7.48$ s to generate $10^{4}$ such waveforms, making it about 60 to 65 times faster than traditional waveform approximation methods. This speed advantage enables rapid parameter estimation and real-time gravitational wave searches. With higher precision, it will support low-latency detection and broader applications in multi-messenger astrophysics.
preprint: APS/123-QED
I Introduction
Since the first direct detection of the binary black hole (BBH) merger GW150914 by LIGO and Virgo in 2015 [1], gravitational wave (GW) astronomy has entered a new era, enabling direct exploration of extreme astrophysical phenomena. With continuous advancements in detector sensitivity, an increasing number of BBH and binary neutron star (BNS) mergers have been observed [2, 3, 4], providing valuable constraints on the equation of state (EoS) of nuclear matter and insights into tidal interactions in neutron stars [2, 5]. BNS mergers serve as natural laboratories for testing general relativity in the strong-field regime and probing high-density nuclear matter [6]. In particular, measurements of tidal deformability impose stringent constraints on the nuclear EoS, shedding light on neutron star structure and ultra-dense matter properties [7, 8]. Additionally, multi-messenger observations, which combine gravitational waves with electromagnetic counterparts, offer an independent method for measuring cosmological parameters, including the Hubble constant [9, 10].
Due to the significance of BNS systems, accurately modeling gravitational waveforms from their mergers is essential for both detection sensitivity and parameter estimation precision. GW searches rely on matched filtering techniques, which require highly accurate waveform templates, while extracting key physical parameters—such as masses, spins, and tidal deformabilities—demands waveform models with high fidelity. However, modeling BNS waveforms remains challenging due to complex physical effects, particularly spin precession and tidal interactions [11, 12, 13]. Waveform modeling has progressed from computationally expensive numerical relativity (NR) simulations [14, 15, 16], which solve Einstein’s equations directly, to more efficient semi-analytical methods. NR simulations provide high-precision waveforms by capturing strong-field and nonlinear effects but are too costly for large-scale parameter-space studies. Post-Newtonian (PN) approximations [17, 18] describe the inspiral phase analytically under weak-field and slow-motion assumptions but lose accuracy near merger. The effective-one-body (EOB) approach [19, 20, 21] improves upon PN by mapping the two-body problem to an effective single-body motion in a modified spacetime. With NR calibrations, EOB models achieve a balance between accuracy and efficiency. Phenomenological models (IMRPhenom) [12, 22, 23, 24] further enhance computational efficiency by fitting frequency-domain waveform templates to extensive datasets. These models enable rapid waveform generation, facilitating large-scale searches. Despite advancements, challenges remain in computational cost, accuracy, and full parameter-space coverage [25, 26, 27], limiting real-time GW detection and precise parameter estimation.
The rapid development of deep learning has introduced an efficient and accurate approach to gravitational waveform modeling [28, 29, 30, 31]. With strong nonlinear fitting capabilities and high computational efficiency [32], deep learning enables high-precision waveform generation at significantly reduced computational cost. George et al. [33] first applied deep learning to BBH waveforms, achieving real-time performance beyond traditional methods. Schmidt et al. [34] used principal component analysis (PCA) with machine learning to reduce the dimensionality of EOB waveforms, improving computational efficiency. Dax et al. [35] further accelerated waveform generation by leveraging the JAX framework, enabling highly efficient computation and real-time inference. Beyond BBH systems, deep learning has shown promise in BNS and extreme mass-ratio inspirals (EMRIs) waveform modeling. Whittaker et al. [36] employed a conditional variational autoencoder (cVAE) to model post-merger signals, capturing uncertainties in the EoS for probabilistic waveform generation. Chua et al. [37] combined reduced-order modeling with deep learning to accelerate waveform generation for EMRIs, reducing the computational cost of fully relativistic waveforms for LISA data analysis by more than four orders of magnitude. These studies indicate that deep learning accelerates waveform generation and generalizes well across high-dimensional parameter spaces, providing an efficient framework for waveform generation.
Despite progress in deep learning-based waveform modeling, most existing models focus on BBH systems or simplified BNS mergers, with precession and tidal effects remaining underexplored. To address this, we propose a Conditional Autoencoder (cAE) framework for rapid BNS waveform generation, with applications in GW data analysis. Our model efficiently generates waveforms conditioned on system parameters $(\Theta)$, including component masses $(m_{1},m_{2})$, spin components $(S_{1x},S_{1y},S_{1z},S_{2x},S_{2y},S_{2z})$, and tidal deformability $(\Lambda_{1},\Lambda_{2})$, while capturing the high-dimensional evolution of GW signals. Trained on a dataset of $3\times 10^{5}$ BNS waveforms from the IMRPhenomXP_NRTidalv2 [38] model, it includes both precession and tidal effects. To enhance learning efficiency, we adopt the amplitude ($A$)-phase ($\Phi$) representation, where the polarization waveforms $h_{+}(t)$ and $h_{\times}(t)$ are expressed in terms of amplitude and phase independently, to reduce data oscillation. The cAE architecture employs a dual-encoder structure, separately encoding physical parameters and waveform data, which are mapped in latent space before reconstruction. By relying solely on forward propagation, cAE achieves exceptional acceleration in large-scale waveform generation. Benchmark tests on an NVIDIA RTX 3090 GPU and an Intel Xeon Silver 4214R CPU show that for generating a single waveform, cAE requires only $0.12$ s, while IMRPhenomPv2_NRTidal and IMRPhenomXP_NRTidalv2 take $0.62$ s and $0.38$ s, respectively, making cAE approximately 5 times faster. For generating $10^{4}$ waveforms in batch, cAE completes the task in $7.48$ s, compared to $497.75$ s for IMRPhenomPv2_NRTidal and $454.51$ s for IMRPhenomXP_NRTidalv2, yielding speed-ups of about 65 and 60 times, respectively. The model’s accuracy is evaluated through waveform overlap calculations, yielding an average mismatch of $3.97\times 10^{-3}$, corresponding to accuracy 99.6%. These results demonstrate that the proposed framework enables efficient, accurate, and scalable BNS waveform generation with precession and tidal effects, making it well-suited for real-time signal detection and parameter estimation.
The structure of the article is as follows: Section 2 describes the waveform representation and the construction of our dataset. Section 3 introduces the fundamental concepts of autoencoders and presents the architecture and hyperparameter settings of our neural network. Section 4 details the model training and validation process. Section 5 evaluates the accuracy and generation efficiency of our model. Finally, Section 6 provides a summary and discusses future research directions.
II Data simulation
This study constructs a dataset of simulated BNS gravitational waveforms to train a cAE. The dataset spans a broad range of physical parameters ($\Theta$), including component masses, spins, and tidal deformability, and provides the corresponding amplitude and phase representations. This formulation enhances the efficiency of deep learning models in capturing waveform structures and their dependencies on $\Theta$.
II.1 Waveform Representation
Gravitational waves are typically characterized by two polarization components, $h_{+}$ and $h_{\times}$, expressed as
$h(t)=h_{+}(t)+ih_{\times}(t).$ (1)
However, directly learning $h_{+}(t)$ and $h_{\times}(t)$ in the time domain is computationally demanding and may hinder training convergence due to waveform complexity. To improve learning efficiency, we adopt an amplitude-phase representation, where $h_{+}$ is treated as the real part and $h_{\times}$ as the imaginary part. The corresponding amplitude $A(t)$ and cumulative phase $\Phi(t)$ are given by
$A(t)=\sqrt{h_{+}^{2}+h_{\times}^{2}},\quad\Phi(t)=\tan^{-1}\left(\frac{h_{\times}}{h_{+}}\right).$ (2)
This representation reduces data dimensionality while enhancing physical interpretability. The amplitude $A(t)$ captures the overall intensity variation of the gravitational wave, while the cumulative phase $\Phi(t)$ describes its temporal evolution, offering a depiction of the underlying dynamics. To further standardize waveform properties, we apply phase normalization,
$\Phi(t)=\Phi(t)-\Phi(t_{0}),$ (3)
which aligns all waveforms to an initial phase of zero. This adjustment improves dataset consistency and stabilizes model training by minimizing phase discrepancies across waveforms, facilitating a more effective learning of parameter dependencies in waveform evolution.
II.2 System Parameter Selection
Previous studies on gravitational waveform modeling have largely focused on BBH systems, while investigations of BNS waveforms remain relatively limited. Moreover, most deep learning models assume either non-spinning neutron stars or perfectly aligned spins, without systematically accounting for precession effects. Tidal deformation, which significantly influences the phase evolution of BNS waveforms, is often simplified using point-mass approximations, leading to the omission of crucial tidal contributions. A further limitation lies in the restricted parameter space of existing datasets. For instance, the mass range is typically constrained to $[1.2,2.0]M_{\odot}$, spin magnitudes remain small ($|S|<0.5$), and the tidal deformability parameters $\Lambda_{1},\Lambda_{2}$ have limited coverage, reducing the ability to capture variations across different EOS. These constraints hinder the generalization capability of deep learning models and limit their applicability in real GW detection. To address these challenges, we develop a machine learning framework that systematically incorporates both precession and tidal effects, enhancing waveform modeling for BNS systems.
To ensure the model effectively learns the key features of BNS waveforms, we select a set of $\Theta$ spanning a broad region of the parameter space, as summarized in TABLE 1. Specifically, we set the luminosity distance to 1 Mpc, and fix both the inclination angle and the coalescence phase to 0.
$\Theta$ Range
$m_{1}$ Uniform$[1,3]M_{\odot}$
$m_{2}$ Uniform$[1,3]M_{\odot}$
$S_{1x}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{1y}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{1z}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{2x}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{2y}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{2z}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$\Lambda_{1}$ Uniform$[0,500]$
$\Lambda_{2}$ Uniform$[0,500]$
Table 1: Range of $\Theta$ for the BNS system in the training dataset ($m_{2}<m_{1}$).
II.3 Construction of Dataset
We construct a dataset of gravitational waveforms for BNS systems to train the cAE. The dataset comprises $3\times 10^{5}$ samples generated using the IMRPhenomXP_NRTidalv2 waveform model, which incorporates both precession and tidal effects. The data generation process consists of parameter sampling, waveform computation, data preprocessing and normalization. $\Theta$ of the BNS system are drawn from uniform distributions, which is shown in TABLE 1. The time-domain GW signals $h_{+}(t)$ and $h_{\times}(t)$ are computed using pycbc.waveform.get_td_waveform[39], followed by trimming to remove leading and trailing zero values. The amplitude $A(t)$ and phase $\Phi(t)$ are then computed using Eq. (2), as shown in FIG. 1.
Figure 1: Input samples in time domain. Top: amplitude curve $A(t)$; middle: phase curve $\Phi(t)$; bottom: waveform strain $h(t)$.
To maintain data uniformity, waveforms are standardized and zero-padded to a fixed duration of $2$ seconds with a sampling rate of 4096 Hz, resulting in $8192$ points per sample. Each waveform segment is extracted from the final $2$ seconds before the ringdown phase, ensuring that the dataset captures the complete merger and ringdown stages while also covering part of the inspiral phase. The 2-second simulated waveform is sufficiently long to encompass the stage of prominent tidal effects preceding the merger [40], enabling the model to learn the characteristic evolution of waveforms across different dynamical regimes.
The dataset consists of three components: $X_{\text{train}}$, which contains the BNS system’s $\Theta$ as input features; $y_{A}$, representing the amplitude data; and $y_{\Phi}$, representing the phase data. During cAE training, we normalize the data to ensure stability and consistency. The $X_{\text{train}}$ are processed using Min-Max Normalization [41], which rescales the data to the range $[0,1]$, ensuring a uniform scale across different parameters and improving training stability:
$X^{\prime}=\frac{X-X_{\min}}{X_{\max}-X_{\min}},$ (4)
where $X$ represents the original data, $X_{\min}$ and $X_{\max}$ are the minimum and maximum values of each parameter, and $X^{\prime}$ is the normalized data. For $y_{A}$ and $y_{\Phi}$, we apply standardization [42], which ensures zero mean and unit variance to eliminate scale differences and enhance training stability:
$X^{\prime}=\frac{X-\mu}{\sigma},$ (5)
where $\mu$ and $\sigma$ denote the mean and standard deviation of the data, respectively. Since the amplitude and phase exhibit distinct evolution patterns, two separate cAEs are trained: one for learning a compact representation and reconstruction of $y_{A}$, and another for modeling $y_{\Phi}$. During training, $X_{\text{train}}$ is provided as a conditional variable to the autoencoder, ensuring that the latent representation $z$ effectively captures the dependence of waveforms on $\Theta$.
III Autoencoder And Conditional Autoencoder
The autoencoder (AE) is an unsupervised learning model widely used for data dimensionality reduction, feature extraction, and generative tasks. It consists of an encoder and a decoder, learning a low-dimensional representation of the data by minimizing reconstruction error. The variational autoencoder (VAE) extends this framework by introducing probabilistic modeling, enforcing a smoother latent variable distribution, which enhances the generative capability. The cAE further incorporates external conditional constraints, enabling the model to generate samples corresponding to specific data distributions based on input conditions, making it particularly relevant for GW waveform modeling.
This section first introduces the fundamental concepts of AE, VAE, and cAE, discussing their applicability to GW waveform generation. Subsequently, we provide a detailed description of the proposed cAE-based waveform generation framework, including the separate cAE architectures designed for phase and amplitude modeling, along with their respective hyperparameter settings.
III.1 Concepts of Autoencoder and Conditional Autoencoder
We employ autoencoders (AEs) [43] to reduce the dimensionality of complex GW waveforms (Amplitude $A$ and Phase $\Phi$) while ensuring accurate reconstruction. As shown in FIG. 2(a), an AE consists of an input $h^{(i)}$ ($A$ or $\Phi$), an encoder $q_{\alpha}(z\mid h)$, a latent variable $z^{(i)}\in\mathbf{R}^{d}$, and a decoder $p_{\beta}({\hat{h}}\mid z)$. The encoder projects the input ($A$ or $\Phi$) into a lower-dimensional latent space $z^{(i)}$, where the dimension $d$ of $z^{(i)}$ can be adjusted based on specific task requirements. The decoder then reconstructs the $\hat{h}^{(i)}$ from the latent representation through an upsampling process. The $\alpha$ and $\beta$ refers to the learned model parameters, such as weights and biases, obtained after training. To measure the similarity between the reconstructed $\hat{h}^{(i)}$ and the target $h^{(i)}$, the AE employs the mean squared error (MSE) as the reconstruction loss:
$L_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}\left\|h^{(i)}-\hat{h}^{(i)}\right\|^{2},$ (6)
where $N$ is the total number of training samples, $h^{(i)}$ represents the target ($A$ or $\Phi$), and $\hat{h}^{(i)}$ is the reconstructed $A$ or $\Phi$. By minimizing $L_{\text{MSE}}$, the model updates its weights and biases, ensuring that $\hat{h}^{(i)}$ closely approximates $h^{(i)}$. Unlike traditional linear methods such as PCA [44], autoencoders (AEs) can capture the nonlinear features of GW signals more effectively [45]. While PCA is efficient for simple signals, its linear projections may miss important features in the nonlinear phases of GW evolution [46, 47]. In contrast, AEs reduce dimensionality through nonlinear mappings, preserving key physical features such as orbital dynamics, tidal effects, and ringdown. This enables better generalization, parameter recovery, and interpolation across the waveform space [48, 49].
Figure 2: (a) Structure of an AE. (b) Structure of a variational VAE.
FIG. 2(b) illustrates the structure of a VAE [50]. Unlike standard AEs, VAEs introduce probabilistic modeling between the encoder and decoder, ensuring that the latent variable $z^{(i)}$ is not a fixed deterministic value but is instead sampled from a distribution defined by the encoder’s output mean $\mu_{\alpha}(h^{(i)})$ and variance $\sigma_{\alpha}^{2}(h^{(i)})$. Specifically, VAEs utilize the reparameterization trick to obtain latent variables:
$z^{(i)}=\mu_{\alpha}(h^{(i)})+\sigma_{\alpha}(h^{(i)})\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,I).$ (7)
This approach allows gradients to propagate through the sampling process, making it possible to optimize the network using gradient-based methods. The VAE training objective consists of the reconstruction loss and the Kullback-Leibler (KL) divergence loss [51, 52]. The reconstruction loss measures the difference between the decoder’s output $\hat{h}^{(i)}$ and the input waveform $h^{(i)}$, typically computed using the negative log-likelihood:
$L_{\text{recon}}=\mathbb{E}_{q_{\alpha}(z\mid h)}\left[-\log p_{\beta}(\hat{h}\mid z)\right].$ (8)
The KL divergence loss ensures that the learned latent variable distribution $q_{\alpha}(z\mid h)$ approximates a predefined prior distribution, typically a standard normal distribution $p(z)=\mathcal{N}(0,I)$:
$L_{\text{KL}}=D_{\mathrm{KL}}\left(q_{\alpha}(z\mid h)\|p(z)\right).$ (9)
The final VAE objective function is given by:
$L_{\text{VAE}}=L_{\text{recon}}+\kappa L_{\text{KL}},$ (10)
where the hyperparameter $\kappa$ controls the weight of the KL divergence loss, regulating the structure of the latent space.
Although traditional AEs and VAEs perform well in capturing the low-dimensional structure and nonlinear features of data, their generative processes typically rely solely on the data itself. As a result, they lack the capacity to explicitly incorporate known physical priors into the latent representations. In other words, standard AE/VAE models in unsupervised learning tend to capture the dominant variations in the data, but they cannot guarantee that the generated waveforms strictly adhere to physical constraints. To address this limitation and further enhance the physical interpretability and controllability of waveform generation, we introduce the cAE [53]. In the cAE framework, additional physical parameters $\Theta$ (as shown in TABEL 1) are incorporated as conditional inputs and jointly mapped with waveform data into a low-dimensional latent space. In this manner, the cAE not only inherits the advantages of AE/VAE in nonlinear dimensionality reduction and data reconstruction, but also enables the explicit embedding of physical constraints into the latent variables, thereby generating waveforms that better reflect realistic astrophysical properties.
Our cAE employs a dual-encoder structure: one encoder processes the original $h$, while the other encodes the $\Theta$, generating latent representations $z_{A/\Phi}$ and $z_{\Theta}$, respectively. These are then mapped in the decoder to reconstruct the waveform $\hat{h}$. The total loss of training cAE consists of a reconstruction loss and a latent consistency loss:
$L_{\mathrm{cAE}}=\frac{1}{N}\sum_{i=1}^{N}\bigl{\lVert}h_{i}-\hat{h}_{i}\bigr{\rVert}_{\text{MAE}}+\lambda\frac{1}{N}\sum_{i=1}^{N}\bigl{\lVert}z_{A/\Phi,i}-z_{\Theta,i}\bigr{\rVert}_{\text{MSE}}^{2}.$ (11)
The first term measures the difference between the reconstructed waveform $\hat{h}_{i}$ and the target waveform $h_{i}$ using Mean Absolute Error (MAE), while the second term ensures consistency between the latent representations $z_{A/\Phi,i}$ and $z_{\Theta,i}$ using Mean Squared Error (MSE). The hyperparameter $\lambda$ balances the two losses, in our study, we fixed it as 1. In the following, we provide a detailed description of our model architecture.
III.2 Architecture and Hyperparameters of the Model
Our study employs the cAEs to model the amplitude and phase of GW waveforms from BNS mergers. The model consists of two independent cAE, each responsible for learning a low-dimensional representation of either the amplitude $y_{A}$ or phase $y_{\Phi}$ and reconstructing waveforms conditioned on $\Theta$. Each cAE comprises two encoders (Encoder 1 and Encoder 2) and a decoder. Encoder 1 processes the amplitude or phase data, while Encoder 2 encodes the $\Theta$, and the two latent representations are combined in the latent space before being mapped back to a complete waveform by the decoder. The overall model architecture is shown in FIG. 3, where Training and Test sections correspond to the training and inference workflows, while the remaining sections detail the structural components.
Figure 3: Training and testing framework of the our model, along with detailed structures of individual components, the structure of the Transformer and ResNet block are shown in Appendix A (FIG 10).
As shown in FIG. 3, our framework consists of two main branches: one for waveform encoding and reconstruction, and one for encoding $\Theta$. The goal is to align latent representations from both branches while ensuring accurate waveform reconstruction.
Encoder 1 is used independently for amplitude and phase inputs. Each encoder processes a normalized 1D data through a Conv1D layer with stride 2, followed by three ResNet blocks for local feature extraction and two Transformer blocks for capturing long-range dependencies. A global average pooling layer compresses the temporal dimension, and a final dense layer maps the features into a latent space ($z_{A}$ or $z_{\Phi}$).
Encoder 2 takes $\Theta$ (TABLE 1) as input. These are passed through four fully connected layers, followed by two Transformer blocks. The output is projected into the same latent space as Encoder 1, producing $z_{\Theta}$.
Decoder takes the latent variable from Encoder 1 and reconstructs the waveform. It expands the latent dimension via a dense layer, then applies two Transformer blocks and three upsampling ResNet blocks using transposed Conv1D layers, recovering the waveform shape.
Latent variables and loss functions: During training, both the waveform and the $\Theta$ are encoded into their respective latent representations, denoted as $z_{A/\Phi}$ and $z_{\Theta}$. To ensure that the decoder can accurately reconstruct the input waveform and that both latent spaces are aligned, we define two loss components: First, the reconstruction loss is computed as the MAE between the input waveform $x$ and the reconstructed waveform $\hat{h}$:
$L_{\text{rec}}=\frac{1}{N}\sum_{i=1}^{N}\left\|h^{(i)}-\hat{h}^{(i)}\right\|_{\text{MAE}}.$ (12)
Second, the latent consistency loss penalizes the difference between the latent vector produced by Encoder 1 ($A/\Phi$) and Encoder 2 ($\Theta$):
$L_{\text{latent}}=\frac{1}{N}\sum_{i=1}^{N}\left\|z_{{A/\Phi}}^{(i)}-z_{\Theta}^{(i)}\right\|_{\text{MSE}}^{2}.$ (13)
The total loss is then defined as a weighted sum of the two as Eq. (11):
$L_{\text{total}}=L_{\text{rec}}+\lambda L_{\text{latent}},$ (14)
where $\lambda$ is a balancing coefficient. In our study, we set $\lambda=1$ to equally weight reconstruction accuracy and latent alignment.
With the model architecture defined, we describe the training procedure and validation setup used to evaluate the model’s performance in the following section.
IV Training and Verification
Our model is trained using the Adam optimizer with an initial learning rate of $10^{-4}$. To enhance training stability and convergence speed, the following learning rate scheduling mechanism is implemented: if the validation loss (val_loss) does not improve for 7 consecutive epochs, the learning rate is reduced by a factor of 0.7, with a minimum threshold of $10^{-8}$. Additionally, an early stopping strategy is employed to prevent overfitting—if the validation loss shows no significant improvement within 15 epochs, training is automatically halted, and the model parameters are reverted to those corresponding to the best validation performance. The training is conducted on a single NVIDIA RTX 3090 GPU, with Amplitude ($A$) and phase ($\Phi$) models trained separately. The total training process took approximately one week on a single RTX 3090. The dataset consists of $3\times 10^{5}$ samples, with 90% used for training and 10% for validation. The batch size is set to 10, and the maximum training epochs are limited to 500. The training of both the amplitude and phase models took approximately a total of seven days. FIG. 4 illustrates the training loss evolution for both the amplitude and phase models.
Figure 4: Training and validation loss curves for the amplitude and phase models.
To evaluate the accuracy of cAE-generated waveforms, we compute the mismatch between the model-generated waveforms and the IMRPhenomXP_NRTidalv2 target waveforms. The evaluation is based on two key metrics: Overlap [54, 55, 56] and Mismatch, which quantify the waveform reconstruction quality. The analysis is conducted in the frequency domain after applying a Fourier transform to the time-domain waveforms, allowing for a more effective comparison of waveform similarities. In GW data analysis, the inner product of two waveforms is typically defined as a noise-weighted integral over the frequency domain, incorporating the power spectral density $S_{n}(f)$:
[…]
where $\tilde{h}_{1}(f)$ and $\tilde{h}_{2}(f)$ are the Fourier transforms of $h_{1}(t)$ and $h_{2}(t)$, respectively, and $S_{n}(f)$ represents the power spectral density (PSD) of the detector noise. This weighted inner product provides a measure of how well two waveforms match in the presence of detector noise. Since our task does not require a realistic noise model, we set $S_{n}(f)=1$.
To eliminate the influence of normalization, each waveform is rescaled to satisfy the unit-norm condition:
$\hat{h}(t)=\frac{h(t)}{\sqrt{\langle h\mid h\rangle}}.$ (16)
The Overlap between two waveforms is then computed by maximizing the inner product over different time shifts $t_{c}$ and phase shifts $\phi_{c}$:
$O(h_{1},h_{2})=\max_{t_{c},\,\phi_{c}}\,\frac{\langle\hat{h}_{1}\mid\hat{h}_{2}\rangle}{\sqrt{\langle\hat{h}_{1}\mid\hat{h}_{1}\rangle\,\langle\hat{h}_{2}\mid\hat{h}_{2}\rangle}}.$ (17)
Based on this Overlap metric, the Mismatch is defined as:
$M(h_{1},h_{2})=1-O(h_{1},h_{2}).$ (18)
Here, $M(h_{1},h_{2})$ quantifies the dissimilarity between the two waveforms, where lower values indicate a higher similarity between the model-generated waveforms and the target physical waveforms.
V Results and Analysis
In this section, we present the accuracy and efficiency of the model on the test set. Furthermore, we analyze various factors that influence the precision of the generated waveforms.
V.1 Mismatch Evaluation
To assess the accuracy of cAE-generated waveforms, we regenerated $3\times 10^{5}$ test samples based on the $\Theta$ ranges listed in TABLE 1 and computed their mismatch values .
Figure 5: The mismatch distribution of the test sample.
FIG. 5 presents the mismatch distribution for $h_{+}$ and $h_{\times}$, showing that their averaged mismatch values are $3.9717\times 10^{-3}$ and $3.9770\times 10^{-3}$, respectively, corresponding to an average waveform accuracy of 99.6%, considering both precession and tidal effects in BNS. Some more detailed charts illustrating the reconstruction results for samples with different orbital cycle numbers are provided in FIG. 11-FIG. 14 in the Appendix A.
To further analyze the factors influencing mismatch, we computed the orbital cycle number for each sample. The time-domain signal $s(t)$ undergoes a Hilbert transform to obtain its analytic representation[57]:
$s_{a}(t)=s(t)+i\,\mathcal{H}\{s(t)\},$ (19)
where ${}\mathcal{H}\{\cdot\}$ denotes the Hilbert transform. The instantaneous phase is then extracted from the analytic signal:
$\phi(t)=\arg\bigl{(}s_{a}(t)\bigr{)}.$ (20)
To eliminate phase discontinuities, we apply phase unwrapping to obtain a monotonic phase function $\tilde{\phi}(t)$ and compute the total phase difference:
$\Delta\phi=\tilde{\phi}(T)-\tilde{\phi}(0).$ (21)
Finally, the orbital cycle number is given by:
$\text{NumCycles}=\frac{\Delta\phi}{2\pi}.$ (22)
FIG. 6(a) shows the test set’s cycle number distribution with a averaged cycle number of 265 and 68% of the samples falling within the range of 231 to 307 cycles.
Figure 6: (a) The test set’s cycle number distribution. (b) Relationship between the number of cycles and mismatch (log10) in the test set. The color bar indicates the number of samples in each hexagonal bin.
As shown in FIG. 6(b), most samples cluster around 250 cycles, where the mismatch remains relatively low. However, when the cycle number exceeds 300, the mismatch begins to increase. This trend, consistent with FIG. 6(b), indicates that an increasing cycle number significantly impacts the mismatch, particularly when exceeding 300, where the model’s reconstruction accuracy deteriorates.
Figure 7: (a) Relationship between binary mass and cycle number in the test set. (b) Relationship between binary mass, cycle number, and mismatch. (c) Relationship between tidal deformability and mismatch in the test set. (d) Relationship between spin parameters and mismatch.
As shown in FIG. 7(a), lower-mass binaries exhibit more waveform cycles, whereas increasing mass results in fewer cycles, so that mismatch values are relatively high in both low-mass (high-cycle) and high-mass (low-cycle) regions (as shown in FIG. 7(b)). This trend is primarily due to two factors: (1) as the number of waveform cycles increases, the time-domain complexity of the signal rises, making it more challenging for the network to learn; (2) high-cycle samples are underrepresented in the training set, reducing the model’s generalization ability in this region. When the cycle number falls between 200 and 250, the mismatch is approximately $3.1\times 10^{-3}$, increasing slightly to $3.2\times 10^{-3}$ in the 250–300 range. However, a significant rise to $7.2\times 10^{-3}$ is observed for 300–350 cycles, and once the cycle number exceeds 350, the mismatch escalates sharply, reaching $4.6\times 10^{-2}$ in the 350–400 range.
In order to investigate the impact of tidal deformation and spin parameters on mismatch, we plotted the mismatch distributions under different tidal deformability and total spin conditions in FIG. 7. FIG. 7(c) shows that the mismatch distribution under different tidal deformabilities is random, which suggests that, under the conditions of this study, tidal deformability has a negligible impact on waveform generation accuracy. Similarly, FIG. 7(d) indicates that spin parameters also exhibit a weak correlation with mismatch, implying that precession effects do not significantly degrade the model’s waveform accuracy. This observation is further supported by FIG. 15 in the Appendix A, where no strong mismatch trend is observed in the joint distributions of spin and tidal deformability parameters.
Figure 8: Relationship between average mismatch and total mass in the test set.
Finally, we analyze the relationship between total mass and mismatch, as shown in FIG. 8. The results indicate that samples with total masses between $3.0M_{\odot}$ and $5.0M_{\odot}$ exhibit the lowest average mismatch. That is consistent with FIG. 7(b), further confirming that our model achieves the highest accuracy when generating waveforms with total binary masses between $3.0M_{\odot}$ and $5.0M_{\odot}$ and cycles between 200 and 250. Additionally, the joint distribution of binary mass in FIG. 15 of the Appendix A reveals a clear trend of increasing mismatch in both high-mass and low-mass regions.
V.2 Waveform Generation Efficiency
To evaluate the efficiency of waveform generation, we compared the cAE model with several traditional waveform approximation methods by measuring the total computation time for different batch size. Specifically, we selected four widely used models: SpinTaylorT1, IMRPhenomPv2_NRTidal, IMRPhenomPv2_NRTidalv2, and IMRPhenomXP_NRTidalv2. All experiments were conducted using an NVIDIA RTX 3090 GPU and an Intel(R) Xeon(R) Silver 4214R CPU, with batch size ranging from 1 to 10,000. The results indicate that the cAE significantly outperforms traditional methods in computational efficiency, particularly for large-scale batch waveform generation.
For generating a single waveform, the cAE requires $0.12$ s, which is shorter than SpinTaylorT1 ($0.29$ s), IMRPhenomPv2_NRTidalv2 [38] ($0.39$ s), IMRPhenomXP_NRTidalv2 ($0.38$ s), and IMRPhenomPv2_NRTidal [58] ($0.62$ s). For the large-scale generation of 10,000 waveforms, the cAE completes the task in only $7.48$ s, while SpinTaylorT1, IMRPhenomXP_NRTidalv2, IMRPhenomPv2_NRTidalv2, and IMRPhenomPv2_NRTidal require $486.86$, $454.51$, $462.87$, and $497.75$ s, respectively. This indicates that our model can efficiently accelerate waveform generation when generating large batch size. A detailed comparison of computation times is presented in FIG. 9 and TABLE 2. Further analysis of total computation time across varying batch size reveals that the cAE exhibits near-sublinear scaling behavior. As the batch size increases from $1$ to $10$, $100$, $500$, $1{,}000$, and $10{,}000$, the total generation time modestly rises from $0.12$ s to $0.12$, $0.20$, $0.50$, $0.86$, and $7.48$ s, respectively. This scaling highlights the excellent efficiency and scalability of the cAE model for batch processing.
In contrast, traditional waveform approximation methods exhibit significantly steeper increases in total computation time as batch size grow. Specifically, SpinTaylorT1 computation time rises from $0.29$ s (batch size = 1) to $486.86$ s (batch size = 10,000). IMRPhenomPv2_NRTidal grows from $0.62$ s to $497.75$ s, IMRPhenomPv2_NRTidalv2 from $0.39$ s to $462.87$ s, and IMRPhenomXP_NRTidalv2 from $0.38$ s to $454.51$ s. These results underscore that without multi-core parallel optimization, traditional methods become notably less efficient compared to the cAE for large-scale waveform generation.
This advantage makes our model particularly suitable for real-time waveform generation and large-scale parameter space sampling, providing a new and feasible approach for efficient GW data analysis.
Figure 9: Comparison of computation time between cAE and traditional waveform approximation. The horizontal axis represents the number of generated samples, while the vertical axis represents the computation time.
Table 2: Waveform generation time (s) for different batch size using the cAE and some waveform approximation methods.
Approximants Batch size 1 10 50 100 500 1000 5000 10000
cAE 0.1188 0.1229 0.1895 0.1954 0.5030 0.8595 3.7589 7.4778
SpinTaylorT1 0.2926 0.8191 3.1774 5.7059 26.0688 51.3171 245.5291 486.8640
IMRPhenomPv2_NRTidal 0.6213 1.5070 3.8230 6.7868 28.2600 53.4714 251.7113 497.7530
IMRPhenomPv2_NRTidalv2 0.3925 0.9189 3.2646 5.6403 23.8314 47.2493 230.3046 462.8673
IMRPhenomXP_NRTidalv2 0.3823 0.9026 3.2442 5.5723 23.6224 46.1181 227.1048 454.5058
VI Summary and Discussion
This study presents an efficient gravitational waveform generation method based on a cAE and applies it to amplitude-phase modeling of BNS systems. Compared to traditional waveform approximation methods such as IMRPhenomXP_NRTidalv2, cAE significantly improves computational efficiency while maintaining high reconstruction accuracy. On a large-scale test dataset, the averaged waveform mismatch is $3.97\times 10^{-3}$, corresponding to an average accuracy exceeding $99.6\%$. Even with precession and tidal effects, cAE maintains high precision across different $\Theta$ ranges, particularly in the high-cycle orbital evolution phase. In terms of efficiency, the cAE model demonstrates a significant advantage over traditional waveform approximation methods. For a single waveform, cAE requires only $0.12$ seconds, compared to $0.38$ seconds for IMRPhenomXP_NRTidalv2 and $0.62$ seconds for IMRPhenomPv2_NRTidal, yielding speedups of approximately $3$ to $5$ times. As the batch size increases, this advantage becomes more pronounced. For 1,000 waveforms, cAE completes the task in $0.86$ seconds, while IMRPhenomXP_NRTidalv2 and IMRPhenomPv2_NRTidal require $46.12$ and $53.47$ seconds, respectively—corresponding to speedups of around $54$ and $62$ times. When generating 10,000 waveforms, cAE requires only $7.48$ seconds, compared to $454.51$ seconds for IMRPhenomXP_NRTidalv2 and $497.75$ seconds for IMRPhenomPv2_NRTidal, achieving speedups of over $60$ times. These results suggest that cAE scales more efficiently with batch size compared to traditional methods. These results suggest that our cAE model is more efficient than traditional waveform approximation methods for generating waveforms at large batch size.
Additionally, this study constructs a large-scale BNS waveform dataset covering an extended physical parameter space, including mass, spin, and tidal deformation, allowing the model to learn waveform characteristics under complex physical conditions. To improve time-series modeling, the cAE architecture integrates ResNet and Transformer mechanisms, enhancing its ability to represent long-duration waveforms and reducing reconstruction errors in high-cycle orbital phases. This design improves generalization and offers new insights for high-precision waveform modeling.
Despite the progress achieved in this study, several aspects require further optimization. The analysis indicates a slight increase in mismatch for both high- and low-mass binaries. In high-mass systems, the short merger time limits the network’s ability to capture waveform details, while the relatively small number of training samples may constrain generalization. Data augmentation or adaptive sampling could mitigate these effects. For low-mass binaries, the higher number of orbital cycles presents additional challenges for accurate waveform fitting, suggesting room for improvement in modeling long-term evolution. The model’s performance under different spin and tidal deformation parameters shows less variation than expected, indicating that sensitivity to extreme parameter regions warrants further investigation. Future work will systematically analyze key factors influencing waveform generation to enhance physical consistency and generalization. While cAE significantly improves computational efficiency over traditional methods, further acceleration strategies, such as TensorRT[59, 60] or ONNX Runtime[61], could further reduce inference time, improving suitability for real-time GW analysis.
Several directions for future research emerge from this study. The impact of latent variable dimensionality on waveform accuracy and efficiency remains unexplored and will be systematically analyzed. Expanding the parameter space to include more complex sources, such as eccentric binaries and a broader range of spin and tidal effects, will further improve applicability. On the architectural side, advanced deep learning methods, including diffusion models [62] and Transformer variants [63], could enhance waveform generation accuracy and efficiency. Moreover, integrating deep learning with JAX-based acceleration techniques [64] may further improve computational efficiency. Additionally, Bayesian inference with neural networks for waveform generation could accelerate parameter estimation [65] for GW events.
In conclusion, the cAE-based waveform generation method proposed in this study offers an efficient and accurate approach to BNS waveform modeling. It shows strong potential for real-time data analysis, large-scale parameter estimation, and GW event identification. As deep learning continues to advance, data-driven methods are expected to play a growing role in GW astronomy, providing more precise and computationally efficient tools for signal modeling and fundamental physics research.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203004), the National Natural Science Foundation of China (Grant No. 12347101), and the Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-MSX0103).
Appendix A
Figure 10: Architectures of the Transformer and ResNet blocks used in our model.
[…]
Figure 15: Corner plot of $3\times 10^{5}$ test set samples, showing the joint distribution of the 10-dimensional parameter space. The diagonal subplots display 1D histograms of each parameter, while the off-diagonal subplots show 2D joint distributions, where the color scale represents the corresponding mismatch value of each sample.